# Thread: A Logic-Based Data Organization Paradigm for How-To Question Answering with Retrieval Augmented Generation

### v1

@@ D.2 Code Template Instruction
Table 10: Instruction that extracts code template and default parameters from the source code.
D.3 Logic Unit Selection Instruction
Table 11 shows the instruction we use to select the most suitable logic unit from a list of candidates.
[System] You are a helpful assistant that selects the most relevant element from <LU_LIST> based on the user’s query in <QUERY> and chat history in <CHAT_HISTORY>. Please respond with the JSON format. Each element in <LU_LIST> is in json format and contains the following fields: { "#type#": "the type of the element, select from the following types: Terminology, FAQ, Step, and Appendix.", "#meta data#": "the description of the troubleshooting guide.", "#prerequisite#": "The prerequisite of this step, before taking the current step, the prerequisites should be finished.", "#header#": "The information describes the intent of the <INFO>.", "#body#": "The action is the content which troubleshoots the incident or explain the #header#. the action may contain code blocks in markdown format, and parameters are replaced with placeholders", "#linker#": "the expected output after taking the #action#. It is defined in the following format in markdown: -If **condition**, then **should_do**. It can contain multiple if-then cases.", "#default_parameters#": "the default parameters that could fill in placeholders in code blocks in #body#." } - The elements in <LU_LIST> contain possible information that can answer the user’s query in <QUERY>. However, they may not be all relevant to the query or useful to answer the user’s query. You should select the most relevant element from the <LU_LIST> based on the user’s query in <QUERY>. - In particular, you should focus on the following fields in the element: #header#, #body#. Most importantly, the <QUERY> needs to match with the #intent# and the #body# has to provide actions to reach the goal of the <QUERY>, please ignore the #linker# and do not map the <QUERY> with #linker#. - As you choosing from <LU_LIST>, you need to check if all the #prerequisite# are met in previous history. If the #prerequisite# is not finished, then it should not be chosen. - Try to select only one element from <LU_LIST>. If it is not possible to select only one element, you can select multiple elements from <LU_LIST>: [ { "INDEX": the index of the element in <LU_LIST>. "INTENT": the #header# of the element, the index starts from 0. "EXPLANATION": justify why you select this node. } ] - If there is no element in <LU_LIST> that can answer the user’s query in <QUERY>, you should try to select the most relevant element to the user’s query considering that the user might use the wrong terminology: [ { "INDEX": the index of the element in <LU_LIST>. "INTENT": the #header# of the element, the index starts from 0. "REPHRASED_QUERY": the rephrased query that you think the user is asking about. "EXPLANATION": justify why you select this node. } ] - Unless you are confident that there is no element in <LU_LIST> that is even close to the user’s query: { "NO_INFO_EXPLANATION": where you give your explanation. } - Your answer should be in the JSON format in a list after <RESPONSE>. [User] <LU_LIST>: {LU_LIST} <QUERY>: {QUERY} <CHAT_HISTORY>: {CHAT_HISTORY}
@@ D.3 Logic Unit Selection Instruction
Table 11: Instruction that selects the most relevant logic unit based on user query and chat history.
Appendix E Details about Mind2web
@@
E.1 Document Generation Instruction
Table 12 shows the instruction we use to generate a specific format of document for Mind2Web dataset.
[System] You are adept at performing website navigation tasks, and you will be provided with simulation data from Mind2Web, designed for developing and evaluating generalist agents capable of following language instructions to complete complex tasks on any website. The data includes a step-by-step execution process, each step encompassing HTML code, Tasks, Previous Actions, and the Element and Action of this step. Note that the Element comes from the HTML code, and if the correct action is not present on the current page, the Element is None, and you should retrieve it from next step’s Previous Actions. Now your task is to write a comprehensive and adaptable reference document that outlines the general process for completing tasks like the given task. This document should serve as a guide for others to perform similar tasks on the same website in the future. So it should not be limited but can use this data to be the example, and should be general enough. Please return the complete reference document that adheres to these guidelines. [User] The format of the documents should be as follows: {FORMAT} The given execution process is as follows: {EXAMPLE}
Table 12: Instruction that generates specific format of document for Mind2web dataset.
E.2 Different formats of Document
Table 13 lists the formats we brainstormed with LLMs, and we use these formats to generate the documents for retrieval in Mind2Web dataset.
@@ E.2 Different formats of Document
Table 13: The description of different formats we use in our experiment.
E.3 Ablations Study of Chat History
Table 14 reveals that incorporating historical steps enhances the performance of LLMs, with an improvement of 5.83 % in Ele. Acc, suggesting that previous steps can help the system make more accurate decisions. So we incorporate historical steps in all our experiments.
Mind2Web Cross-Task
Paradigm Ele. Acc Op. F1 Step SR
GPT-3.5 40.69 49.66 33.91
$w/o.$ historical steps 36.25 45.00 31.38
GPT-4 62.80 60.37 51.81
$w/o.$ historical steps 56.97 59.09 50.38
Table 14: Ablation study of historical steps on Mind2web.
E.4 Ablations Study of Multi-Agent System
We explore the impact of more complex RAG systems on their performance. We adapt the multi-agent system proposed by Wu et al. (2023); An et al. (2024), which is more intricate, to execute Mind2web tasks. In the multi-agent system, multiple LLM-based agents such as the Selector, Action Planner, and Query Generator collaborate, enhancing performance on complex how-to questions. As evidenced in Table 15, there is a noticeable improvement in the overall performance, with the Success Rate on Mind2Web tasks rising by 3.36%. This increment suggests that the deployment of multi-agent systems can substantially improve the capability to manage tasks of greater complexity.
Mind2Web Cross-Task
#Agent Ele. Acc Op. F1 Step SR SR
Single 68.29 69.53 61.94 12.30
Multiple 68.40 71.50 62.18 15.66
Table 15: Analysis of systems with different agents for Dynamic How-To Questions.
Appendix F Details about Wikihow
Table 16 shows the instructions about how to evaluate the result of Wikihow, including extraction of action items and assessment of success.
[System] You are a helpful and precise assistant for checking the quality of the answer. We would like to invite you to evaluate the performance of the system in answering a user’s question in <Question>. I will give you the answer generated by the system in <Generation> and the ground truth answer in <Ground Truth> respectively. Your evaluation will contain five sub-evaluation tasks: 1. Both two answers contain a list of steps. Your task is to extract action items from the provided steps in both answers. The action item is defined as a combination of action and element. Compare the action items to identify similarities. Output the similar action items. Count the count of similar action items. - Your answer should contain the extracted two action item sets (in the format as a list of strings). - Your answer should contain a set of similar action items (in the format of a list of strings). Similar action items are those sharing similar intent or achieving similar goals. Each similar action pair in the list should be in the format of "similar action item from action item set1 / similar action item from action item set2" - Your answer should contain the count of similar action items. 2. Can <Generation> completely solve the user’s question? - Your answer should be "Yes" or "No". - Your answer should contain the reason(s) for your choice. You should not focus on the length of the answer or the details of the answer, but you should focus on whether the steps could solve the user’s question and the quality of the steps compared with the ground truth. Your output should be in the following format in JSON: { "Subtask1": { "Action items in Generation": ["action item 1", "action item 2", ...], "Action items in Ground Truth: ["action item 1", "action item 2", ...], "Similar action items": ["similar action item 1", "similar action item 2", ...], "Count of similar action items": 2 }, "Subtask2": { "Choice": "Yes" or "No", "Reason": "reason for your choice" } } [User] Here is the user’s question <Question>: {Question} The answer from system <Generation> is: {Generation} The ground truth answer <Ground Truth> is: {Ground Truth}
@@ Appendix F Details about Wikihow
Table 16: Instruction that evaluates the generated answer compared with ground truth for Wikihow.
Appendix G Details about RAG System
We take the scenario of Mind2web to show the instructions we used in our RAG system.
G.1 Instruction of Baselines
Table 17 shows the instruction we use in our RAG system of baselines.
[System] You are a helpful assistant who is great at website design, navigation, and executing tasks for the user. Now please proceed with the <CURRENT_STEP> and make your choice, remember that only based on the helpful document information from <DOC_CONTEXT> and the previous step chat history between user and assistant in <CHAT_HISTORY>. Your response should be in the format of "Answer: C. Action: SELECT Value: Pickup". The answer is A, B, C…, the Action comes from [CLICK, TYPE, SELECT] and the Value is not always needed. [User] <DOC_CONTEXT>: {DOC_CONTEXT} <CHAT_HISTORY>: {CHAT_HISTORY} <CURRENT_STEP>: {CURRENT_STEP}
Table 17: Instruction that is used for the baselines of RAG system on Mind2web.
G.2 Instruction of Thread Paradigm
Table 18 shows the instruction we use in our RAG system utilizing the Thread paradigm, which is different from baselines as it requires pointing out the query for next-time retrieval.
[System] You are a helpful assistant who is great at website design, navigation, and executing tasks for the user. Now please proceed with the <CURRENT_STEP> and make your choice, remember that only based on the helpful structured document information from <LU>, and the previous step chat history between user and assistant in <CHAT_HISTORY>. Your response should be in the format of JSON: { "CHOICE": the choice you make from A B C ..., "ACTION": the corresponding action choosing from [’CLICK’, TYPE’, SELECT’], "VALUE": the corresponding value if needed, "INTENT": the intent of the next step, which should be retrieved and judged from the "if" conditions in #output# from <LU> according to the current step and actions and choose the corresponding "then" outcome, do not guess it based on current Task in <CURRENT_STEP> by yourself unless the <LU> is irrelevant to <CURRENT_STEP>, } [User] <LU>: {LU} <CHAT_HISTORY>: {CHAT_HISTORY} <CURRENT_STEP>: {CURRENT_STEP}
@@ G.2 Instruction of Thread Paradigm
Table 18: Instruction that is used for the RAG system utilizing Thread Paradigm on Mind2Web.

### v2

@@ Appendix B Example of Dataset
Instruction that extracts code template and default parameters from the source code.
Each element in <LU_LIST> is in json format and contains the following fields:
"#type#": "the type of the element, select from the following types: Terminology, FAQ, Step, and Appendix.",
"#meta data#": "the description of the troubleshooting guide.",
"#prerequisite#": "The prerequisite of this step, before taking the current step, the prerequisites should be finished.",
"#header#": "The information describes the intent of the <INFO>.",
"#body#": "The action is the content which troubleshoots the incident or explain the #header#. the action may contain code blocks in markdown format, and parameters are replaced with placeholders",
"#linker#": "the expected output after taking the #action#. It is defined in the following format in markdown: -If **condition**, then **should_do**. It can contain multiple if-then cases.",
"#default_parameters#": "the default parameters that could fill in placeholders in code blocks in #body#."
- The elements in <LU_LIST> contain possible information that can answer the user’s query in <QUERY>. However, they may not be all relevant to the query or useful to answer the user’s query. You should select the most relevant element from the <LU_LIST> based on the user’s query in <QUERY>.
- In particular, you should focus on the following fields in the element: #header#, #body#. Most importantly, the <QUERY> needs to match with the #intent# and the #body# has to provide actions to reach the goal of the <QUERY>, please ignore the #linker# and do not map the <QUERY> with #linker#.
- As you choosing from <LU_LIST>, you need to check if all the #prerequisite# are met in previous history. If the #prerequisite# is not finished, then it should not be chosen.
- Try to select only one element from <LU_LIST>. If it is not possible to select only one element, you can select multiple elements from <LU_LIST>:
"INDEX": the index of the element in <LU_LIST>.
"INTENT": the #header# of the element, the index starts from 0.
"EXPLANATION": justify why you select this node.
- If there is no element in <LU_LIST> that can answer the user’s query in <QUERY>, you should try to select the most relevant element to the user’s query considering that the user might use the wrong terminology:
"INDEX": the index of the element in <LU_LIST>.
"INTENT": the #header# of the element, the index starts from 0.
"REPHRASED_QUERY": the rephrased query that you think the user is asking about.
"EXPLANATION": justify why you select this node.
- Unless you are confident that there is no element in <LU_LIST> that is even close to the user’s query:
"NO_INFO_EXPLANATION": where you give your explanation.
- Your answer should be in the JSON format in a list after <RESPONSE>.
[User]
<LU_LIST>: {LU_LIST}
<QUERY>: {QUERY}
<CHAT_HISTORY>: {CHAT_HISTORY}
@@
Instruction that selects the most relevant logic unit based on user query and chat history.
C.2 Mind2web
@@
The data includes a step-by-step execution process, each step encompassing HTML code, Tasks, Previous Actions, and the Element and Action of this step. Note that the Element comes from the HTML code, and if the correct action is not present on the current page, the Element is None, and you should retrieve it from next step’s Previous Actions.
Now your task is to write a comprehensive and adaptable reference document that outlines the general process for completing tasks like the given task. This document should serve as a guide for others to perform similar tasks on the same website in the future. So it should not be limited but can use this data to be the example, and should be general enough.
Please return the complete reference document that adheres to these guidelines.
[User]
The format of the documents should be as follows: {FORMAT}
The given execution process is as follows: {EXAMPLE}
Instruction that generates specific format of document for Mind2web dataset..
We follow Wang et al. (2024) to brainstorm diverse formats of documents for Mind2web dataset used for retrieval, and the results are listed in Table 10.
@@
Table 10: The description of different formats of documents on Mind2web.
C.3 Wikihow
I will give you the answer generated by the system in <Generation> and the ground truth answer in <Ground Truth> respectively. Your evaluation will contain five sub-evaluation tasks:
1. Both two answers contain a list of steps. Your task is to extract action items from the provided steps in both answers. The action item is defined as a combination of action and element. Compare the action items to identify similarities. Output the similar action items. Count the count of similar action items.
- Your answer should contain the extracted two action item sets (in the format as a list of strings). - Your answer should contain a set of similar action items (in the format of a list of strings). Similar action items are those sharing similar intent or achieving similar goals. Each similar action pair in the list should be in the format of "similar action item from action item set1 / similar action item from action item set2" - Your answer should contain the count of similar action items.
2. Can <Generation> completely solve the user’s question? - Your answer should be "Yes" or "No". - Your answer should contain the reason(s) for your choice. You should not focus on the length of the answer or the details of the answer, but you should focus on whether the steps could solve the user’s question and the quality of the steps compared with the ground truth.
Your output should be in the following format in JSON:
"Subtask1": {
"Action items in Generation": ["action item 1", "action item 2", ...],
"Action items in Ground Truth: ["action item 1", "action item 2", ...],
"Similar action items": ["similar action item 1", "similar action item 2", ...],
"Count of similar action items": 2
"Subtask2": {
"Choice": "Yes" or "No",
"Reason": "reason for your choice"
[User]
Here is the user’s question <Question>: {Question} The answer from system <Generation> is: {Generation} The ground truth answer <Ground Truth> is: {Ground Truth}
@@ C.3 Wikihow
Instruction that evaluates the generated answer compared with ground truth for Wikihow.
Appendix D Details about RAG System
We take the scenario of Mind2web to show the instructions we use in our RAG-based QA system.
Your response should be in the format of "Answer: C. Action: SELECT Value: Pickup". The answer is A, B, C…, the Action comes from [CLICK, TYPE, SELECT] and the Value is not always needed.
[User]
<DOC_CONTEXT>: {DOC_CONTEXT} <CHAT_HISTORY>: {CHAT_HISTORY} <CURRENT_STEP>: {CURRENT_STEP}
Instruction that is used for the baselines of RAG system on Mind2web.
Your response should be in the format of JSON:
"CHOICE": the choice you make from A B C ...,
"ACTION": the corresponding action choosing from [’CLICK’, TYPE’, SELECT’],
"VALUE": the corresponding value if needed,
"INTENT": the intent of the next step, which should be retrieved and judged from the "if" conditions in #output# from <LU> according to the current step and actions and choose the corresponding "then" outcome, do not guess it based on current Task in <CURRENT_STEP> by yourself unless the <LU> is irrelevant to <CURRENT_STEP>,
[User]
<LOGIC_UNIT>: {LOGIC_UNIT} <CHAT_HISTORY>: {CHAT_HISTORY} <CURRENT_STEP>: {CURRENT_STEP}
@@ Appendix D Details about RAG System
Instruction that is used for the RAG system utilizing Thread Paradigm on Mind2Web.
Appendix E Limitations
This work focuses on evaluating the effectiveness of Thread by designing how-to questions in three specific scenarios, covering both linear and dynamic how-to questions. However, there are several limitations that point to future directions. First, as our logic-based knowledge base can coexist with the original chunk-based knowledge base, we do not extend our method to factoid questions, such as multi-hop or long-form questions. Second, the experiments primarily use ChatGPT and LLaMA as the backbones. Future work should include evaluations with other LLMs and alternative retrievers such as Contriever to further validate the effectiveness of Thread. Lastly, while extracting logic units involves an initial cost in terms of LLM usage, this is a one-time process. Once the knowledge base is constructed, it provides significant advantages for industrial applications, particularly in terms of subsequent updates and maintenance.